OpenStudy (anonymous):

The radius of the earth is 6370Km and the radius of mars is 3440Km. If an object weights 200N on earth, what will be the gravity on mars? Mars has a mass of 0.11 times the mass of the earth.

OpenStudy (dumbcow):

hey that formula will work, just plug in the numbers and you will get Fg = 75

OpenStudy (dumbcow):

oh the formula needs radius in meters not km, maybe that will help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Hey dumbcow, I deleted my last question. Can you show the work again, please

OpenStudy (anonymous):

the answer is 75N and 3.7m/s2

OpenStudy (anonymous):

\[g=GM/r^2\]

OpenStudy (dumbcow):

\[F = G \frac{m _{1}m _{2}}{r^{2}}\] \[F = (6.67 X 10^{-11})\frac{(20.41)(0.11)(5.9736 X 10^{24})}{6,370,000^{2}}\]

OpenStudy (anonymous):

what does 20.41 means?

OpenStudy (dumbcow):

mass of the object mg = 200, where g is 9.8 m = 200/9.8 = 20.41

OpenStudy (anonymous):

I multiplied that but I didnt get 75

OpenStudy (dumbcow):

oops i put the wrong radius use radius from mars

OpenStudy (anonymous):

ok let me check it out

OpenStudy (anonymous):

I got 7.55929*10^7 when I replace the radius of mars

OpenStudy (dumbcow):

or 75.5 X 10^6 = 75/10^-6 you forgot to add the 3 zeroes to radius to convert it to meters

OpenStudy (anonymous):

Yeah I got 75 finally

OpenStudy (anonymous):

\[g(earth)/g(mars)=[M(earth)x/r^2(earth)] X [r^2(mars)/M(mars)]\]

OpenStudy (anonymous):

what does that formula means. Does that do the same?

OpenStudy (dumbcow):

Then 75 = mg 75 = 20.41g 3.67 = g

OpenStudy (anonymous):

Thanks dumbcow

OpenStudy (anonymous):

still want to know what does that formula means SmRoberts. Please tell me

OpenStudy (anonymous):

\[g = GM/r^2\] where G is the universal constant of gravitation M is mass of the planet r is the radius of the planet

OpenStudy (anonymous):

did you get 75N?

OpenStudy (anonymous):

yes i did

OpenStudy (anonymous):

We Wm --= ---- ge gm Wm=(200/9.8)3.7=75N

OpenStudy (anonymous):

did u get it emunrra?

OpenStudy (anonymous):

Mm/Me 6.5736x10^23kg/5.976x10^24 gm=---------=------------------------------=3.7m/s answer (rm/re)^2 (3440km/6370km)^2

OpenStudy (anonymous):

yes, I did. Thanks. This procedure is clearer :)

OpenStudy (anonymous):

ok good luck to you..lol

OpenStudy (anonymous):

How do I get gm??? what is the original equation?

OpenStudy (anonymous):

please help me. I dont get it